# Thermodynamic and quantum fluctuations of horizon area

### v2

@@ II Quantum fluctuations of black hole entropy
These canonically conjugate variables were used in particular for the calculations of the transition rate from the black hole to the white hole with the same mass.Volovik2022 This rate is determined by the integral $\int_{C}A(K)dK$ over the tunneling trajectory $C$, which connects black and white holes.
@@
$\left<(\Delta A)^{2}\right>\,\geq\frac{1}{K}\left<A\right>=4G\left<A\right$>\,. (4)
@@ III Thermodynamic fluctuations of de Sitter horizon entropy
Since $S=KA=A/4G$, this gives again the Eq.(4). Note that Eq.(6) does not contain the gravitational coupling $K$. This is in agreement with the observation by Jacobson Jacobson1994 that both $K$ and $S$ are renormalized by the very same quantum fluctuations.
The exact relation between the quantum and thermodynamic fluctuations is typical for the systems with horizons. This can be seen also from the processes of splitting of the black hole into the smaller parts. On one hand, such process is quantum, being determined by the macroscopic quantum tunneling, and on the other hand it is determined by the decrease of entropy after splitting,Wilczek2000 ; Volovik2022 which is manifestation of the thermal fluctuations.

### v3

@@ II Quantum fluctuations of black hole entropy
These canonically conjugate variables were used in particular for the calculations of the transition rate from the black hole to the white hole with the same mass.Volovik2022 This rate is determined by the integral $\int_{C}A(K)dK$ over the tunneling trajectory $C$, which connects black and white holes. The extended first law of black hole thermodynamics, which includes $K$ as a thermodynamic variable, is dS=AdK+(1/T)dMdS=-AdK+(1/T)dM, see Ref. Volovik2022 and also Ref. KimNam2025 .
@@
$\left<(\Delta A)^{2}\right>\,\geq\frac{1}{K}\left<A\right>=4G\hbar\left<A\right$>\,. (4)
Here we restored $\hbar$ (keeping $c=1$) to demonstrate that these are quantum fluctuations of the black hole area.
@@ III Thermodynamic fluctuations of de Sitter horizon entropy
The Eq.(6) does not contain $\hbar$, which demonstrates that it describes thermodynamic fluctuations. However, since $S=A/4G\hbar$, this is equivalent to the Eq.(4) describing quantum fluctuations. The exact relation between the quantum and thermodynamic fluctuations is typical for the systems with horizons. This can be seen also from the processes of splitting of the black hole into the smaller parts. On one hand, such process is quantum, being determined by the macroscopic quantum tunneling, and on the other hand it is determined by the decrease of entropy after splitting,Wilczek2000 ; Volovik2022 which is manifestation of the thermal fluctuations. This is one of many examples where gravity serves as a bridge between thermodynamics and quantum mechanics.Wald1999
@@
Note that Eq.(6) does not contain the gravitational coupling $K$. This is in agreement with the observation by Jacobson Jacobson1994 that both $K$ and $S$ are renormalized by the very same quantum fluctuations.
@@ IV Black hole fluctuations via ensemble of Planck scale ”atoms”
The equation (10) also suggests the following variance of the black hole mass $M$:
$\left<(\Delta M)^{2}\right>=\frac{1}{2}m_{\rm P}^{2}\,.$ (13)
This shows that, unlike the traditional approach (see e.g. Ref. Avramov2024 and references therein), in the toy model the black hole is robust to thermodynamic fluctuations. This is because in the toy model the black hole is the equilibrium state. The Hawking radiation represents the thermal/quantum fluctuation. After the emission of Planckons, the entropy of the black hole decreases, but it is subsequently restored in the natural process of absorption of the emitted Planckons.